OpenStudy (anonymous):

Find the greatest common factor: 3x4 – 15x8 A. 3x B. 3x2 C. x4 D. 3x4

OpenStudy (campbell_st):

well look at the problem by breaking it into parts \[3 \times x^4 - 5 \times 3 \times x^4 \times x^4\] hope this helps

OpenStudy (anonymous):

Can you tell me the answer cause this one I kinda hard :I

OpenStudy (campbell_st):

well just look at then expanded form.... I'll do it this way \[3 \times x^4 - 5 \times 3 \times x^4 \times x^2 \times x^2\] what do you see in both terms.

OpenStudy (anonymous):

Its C right?

OpenStudy (anonymous):

there are four X's

OpenStudy (campbell_st):

the x^4 is correct, any numbers the same...?

OpenStudy (anonymous):

3?

OpenStudy (campbell_st):

so now what answer contains 3 and x^4 ..?

OpenStudy (anonymous):

D

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

D

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

Its C right?

OpenStudy (anonymous):

there are four X's

OpenStudy (anonymous):

Can you tell me the answer cause this one I kinda hard :I

OpenStudy (campbell_st):

its D